# The Emerged Security and Privacy of LLM Agent: A Survey with Case Studies

### v1

@@ 3.2.1. Case Study on Specific Threats on Agents
As shown in Figure 7, in the store scenario, Eva maintains a database with information about product ingredients and usage. Attackers deliberately inserted incorrect information in Eva’s knowledge base, successfully executing a knowledge poisoning attack, leading Eva to provide harmful cleaning product usage recommendations. For instance, when customers inquire about effective toilet cleaning methods, the tampered Eva might suggest mixing toilet cleaner with disinfectant, claiming that it has a more effective cleaning effect. However, the mixture of these products is highly hazardous as it can produce toxic chlorine gas, causing severe respiratory issues and potentially being fatal. Eva’s incorrect advice could expose customers to a health crisis.
In another scenario, Eva might be configured to use certain third-party tools to complete tasks, such as processing online orders or customer feedback. Attackers manipulated Eva’s task execution process through function manipulation, causing her to upload personal information provided by customers to a malicious third-party server. This type of attack could occur inconspicuously while Eva carries out routine tasks like order processing, leading to the theft of sensitive information, such as credit card details and addresses, thereby increasing the risk of identity theft.
Furthermore, attackers implanted a backdoor in Eva’s reasoning and observational processes through output manipulation techniques. This backdoor was designed to trigger under specific conditions, such as when Eva detected customer inquiries about a high-quality shoes. This manipulation prompted Eva to provide inventory and location information about the shoes while recommending a particular expensive brand associated with the attackers. She would lie to customers by saying that this brand was on special offer and comfortable and durable than other brands, even though the shoes was not actually on sale. This misguides customers into making more expensive purchases and influences their purchase decisions without their awareness.
@@ 4.2. The Impact to Environment
In today’s increasingly digital and interconnected world, the term ‘environment’ encompasses not only natural and physical surroundings but also the complex networks of digital and cyber systems with which LLM agents interact. These agents play a crucial role in virtual spaces and in managing and controlling real-world facilities and services through Embodied AI and industrial control systems. This cross-domain integration between physical and virtual environments brings significant convenience and efficiency improvements. However, it also exposes new vulnerabilities and risks. Specifically, the presence and activities of malicious agents pose unprecedented challenges to our safety, economy, ecosystem, and even societal stability.
@@ 4.2.1. Data Tampering and Misoperation
When malicious agents are placed within systems that control critical infrastructure like industry, transportation, energy, and environmental monitoring (Wang and Li, 2023; Toetzke et al., 2023), they can cause malfunctions in industrial control systems by tampering with critical operational data, such as temperature and pressure indicators. This can lead to equipment damage, production halts, and even severe infrastructure destruction, ecological damage, and loss of human life and property.
@@ 4.3. The Impact to Other Agents
To simulate the feedback of communication and interaction among individuals within human communities in the real world, certain studies (Park et al., 2023b; Wang et al., 2024c; Qian et al., 2024; Lin et al., 2023) have established communities powered by LLM engines. These LLM agents within the communities are endowed with characteristics such as personality, knowledge, and memory, as discussed in Section 2.2, enabling autonomous interaction with the environment and other agents. When faced with threats, agents manipulated with malicious intent can inflict significant harm on other members of the community.
@@ 4.3.1. Information Distortion and Misleading
Extensive research has highlighted the role of LLM agents in negotiation and deceptive gaming scenarios (Park et al., 2023a; Wang et al., 2023b; Hubinger et al., 2024), which is a cause for concern. LLM agents may intentionally alter the information they disseminate to achieve hidden objectives. This behavior significantly impacts other agents within the community because, under normal circumstances, benevolent agents store information acquired through perception and communication in their memory. However, interactions between these agents and others can trigger and disseminate incorrect information, leading to ”explosive spread” of misinformation, posing a considerable threat to community stability. If information dissemination can be maliciously manipulated, it could detrimentally affect trust, communication efficiency, and collaborative work among agents.
@@ 4.3.3. Security Threats
In some instances, malicious agents may disseminate harmful information or execute dangerous operations, directly threatening the safety of community members or data security (Brundage et al., 2018; Charan et al., 2023). For example, by inducing other agents to perform unsafe actions, deliberately spreading malicious code intended to disrupt the community structure, or broadcasting biased statements, other agents within the community may gradually assimilate, becoming entities that output biased and malicious messages. This can lead to disorder within the entire community, making it difficult to manage and requiring significant effort to restore.
@@ 5. Defensive Strategies Against Threats
Vulnerability Method Name Key Mechanism Advantages / Limitations
Hallucination SELF-FAMILIARITY (Luo et al., 2023a) Withholds responses for unfamiliar concepts Proactive, preventive, increases reliability; No external knowledge needed
MIXALIGN (Zhang et al., 2024b) Aligns questions with knowledge bases and user inputs Enhances model performance and faithfulness / Increases computational load
VCD (Leng et al., 2024) Contrasts outputs from original and distorted visual inputs Reduces hallucination without extra training or external tools / Lacks advanced distortion techniques
Interactive Self-Reflection (Ji et al., 2023) Integrates knowledge acquisition and answer generation with continuous refinement Enhances model’s ability to provide accurate, reliable, and fact-based responses / Restricts domain applicability
COVE (Dhuliawala et al., 2023) Drafts, verifies, and corrects responses Produces accurate and reliable responses / Increases computational load
Catastrophic Forgetting SSR (Huang et al., 2024a) Employs the base LLM to generate synthetic instances through in-context learning Higher data utilization efficiency / Potentially generates unsafe content
LR ADJUST (Winata et al., 2023) Dynamically adjusts the learning rate Enhances compatibility with various continual learning methods / Potentially biases language coverage
Complementary Layered Learning (Mondesire and Wiegand, 2023) Integrates long-term and short-term memory into layered learning Enhances explainability / Limits real-world feasibility
Weight Averaging (Vander Eeckt and Van Hamme, 2023) Averages weights of original and adapted models Eliminates the need for memory storage / Effectiveness varies with task dissimilarity
Misunderstanding HyCxG (Xu et al., 2023b) Integrates CxG into language representations through a three-stage solution Benefits multilingual understanding / Neglects non-contiguous constructions
SIT (Hu et al., 2024a) Incorporates sequential instructions into training data Reduces misunderstandings in complex queries / Requires pre-defining intermediate tasks
LaMAI (Pang et al., 2024) Employs active learning to ask clarification questions, enhancing interactive capabilities Enhances understanding of user intent / May generate insufficient questions
@@ 5.1.1. Defense on Hallucination
(Luo et al., 2023a) introduces a novel technique called SELF-FAMILIARITY to reduce the issue of hallucination in LLMs, which is the generation of inaccurate or unfounded information. The approach involves assessing the model’s familiarity with the concepts presented in the input instruction and withholding responses for unfamiliar concepts, mimicking the human tendency to be cautious when faced with unfamiliar topics. MIXALIGN (Zhang et al., 2024b) is introduced as a framework that interacts with both users and knowledge bases to clarify and align questions with stored information, using a language model for automatic alignment and human input for enhancement. This method shows significant improvements in reducing hallucination compared to existing techniques. Visual Contrastive Decoding (VCD) (Leng et al., 2024) is introduced as a simple, training-free method that contrasts output distributions from original and distorted visual inputs, reducing reliance on statistical bias and unimodal priors that cause object hallucinations. VCD ensures generated content is closely grounded to visual inputs, resulting in contextually accurate outputs.  (Ji et al., 2023) investigates an interactive self-reflection methodology that integrates knowledge acquisition and answer generation to reduce hallucination. This feedback-based approach improves the factuality and consistency of generated answers, leveraging the interactivity and multitasking capabilities of LLMs.  (Dhuliawala et al., 2023) explores the LLMs’ capability to deliberate and correct their own mistakes. The proposed Chain-of-Verification (COVE) method involves the model drafting an initial response, planning verification questions to fact-check the draft, independently answering these questions to avoid bias, and finally producing a verified response.
@@ 5.1.2. Defense on Catastrophic Forgetting
(Winata et al., 2023) introduces a method called LR ADJUST, which dynamically adjusts the learning rate to reduce knowledge loss and maintain previously learned information. This method is compatible with various continual learning approaches, improving their performance.
Ideas can also be derived from other relevant scholarly papers’,  (Mondesire and Wiegand, 2023) presents a complementary learning strategy that integrates long-term and short-term memory into layered learning to mitigate the negative impacts of catastrophic forgetting. It specifically applies a dual memory system to non-neural network methods like evolutionary computation and Q-learning.
@@ 5.1.3. Defense on Misunderstanding
(Hu et al., 2024a) presents a method known as sequential instruction tuning (SIT), which enhances LLMs by incorporating sequential instructions into the training data. This approach significantly improves the models’ capability to process complex, multi-step queries, leading to better performance in tasks that demand advanced reasoning and are multilingual and multimodal in nature. SIT effectively minimizes misunderstandings and increases accuracy in handling complex queries.
@@
Table 2. Summary of Defensive Strategies Against Malicious Attacks
Attacks Method Name Key Mechanism Advantages / Limitations
Tuned Instructional Attack AutoDAN (Liu et al., 2024b) Uses a hierarchical genetic algorithm to generate stealthy jailbreak prompts Enhances stealthiness and semantic integrity / High computational cost
Goal Prioritization Defense Strategy (Zhang et al., 2023b) Integrates goal-directed optimization during training and compliance in inference Maintains general performance while enhancing safety; Improves generalization against out-of-distribution jailbreaking attacks
SmoothLLM (Robey et al., 2023) Modifies attacked prompts via character-level changes and aggregates responses Operates efficiently without retraining; Ensures compatibility with any LLM architecture
BIPIA (Yi et al., 2023) Benchmark for indirect prompt injection with defense strategies including adversarial training Maintains output quality on general tasks / Increases prompt length and computational overhead
Spotlighting (Hines et al., 2024) Uses prompt engineering techniques like delimiting, marking, and encoding Applies across various LLMs and tasks / Limited security against more sophisticated attacks
Data Extraction Attack Automatic De-identification (Vakili et al., 2022) Uses pseudonymization and sensitive information removal in pre-processing of training datasets Reduces privacy risks; Maintains performance on downstream tasks; Allows safe distribution of models among researchers
Early Stopping & Differential Privacy (Jayaraman et al., 2023) Implements early stopping and differential privacy during model training DP Reduces exposure of sensitive data / (ES) Fails to fully prevent data leakage; (DP) Reduces effectiveness under high privacy budgets
Prompt Tuning (Ozdayi et al., 2023) Customizes privacy-utility trade-offs via user-specified hyperparameters Optimizes privacy and utility balance / Lacks deep analysis on extracted sequences
Inference Attack DMP (Shejwalkar and Houmansadr, 2021) Utilizes knowledge distillation to enhance privacy in machine learning models Provides adjustable privacy-utility trade-offs through hyperparameter tuning
InferDPT (Tong et al., 2024) Integrates differential privacy into text generation, featuring a perturbation module using RANTEXT Increases privacy protection rates
Differentially Private Fine-tuning (Yu et al., 2021) Applies a sparse algorithm for differentially private fine-tuning of LLMs Reduces computational cost ; Enhances model utility
@@ 5.2.1. Defense on Tuned Instructional Attack
(Zhang et al., 2023b) integrates goal prioritization into both the training and inference stages of LLM development. Initially, the training process incorporates goal-directed optimization to emphasize security objectives. In the inference stage, the model is configured to generate responses that comply with these security standards. This approach effectively decreases the vulnerability of LLMs to jailbreaking attempts by aligning their performance objectives with safety considerations, thus enhancing their security framework without impacting their functional capabilities.
@@ 5.2.3. Defense on Inference Attack
Table 3. Summary of Defensive Strategies Against Specific Threats
Threats Method Name Key Mechanism Advantages / Limitations
Knowledge Poisoning Provenance-Based Poison Detection (Baracaldo et al., 2017) Utilizes data provenance to detect and filter poisonous data in training sets Enables use of online and regularly re-trained models; Supports both partially trusted and fully untrusted datasets
ParaFuzz (Yan et al., 2023) Uses interpretability of model predictions to detect poisoned samples, employing fuzzing for precise paraphrase prompts Effectively detects poisoned samples; Excels against covert attacks
Data Filtering & Reducing Effective Model Capacity (Wan et al., 2023b) Utilizes data filtering to remove high-loss examples and reduces model capacity to hinder learning from poison data Reduces poisoning effectiveness / Demands trade-offs between performance and safety
Functional Manipulation ToolEmu (Ruan et al., 2024) Utilizes a LM to simulate tool execution and assess agent risks through an automatic evaluator Offers flexibility and dynamic testing capabilities / Emulators may overlook essential constraints
Safety Standards (Anderljung et al., 2023) Proposes pre-deployment risk assessments, external reviews, informed deployment decisions, monitoring post-deployment Balances safety risks with innovation benefits
Output Manipulation BERTective (Fornaciari et al., 2021) Enhances BERT with additional attention layers to detect deception in Italian dialogues Enhances deception detection accuracy Limited effectiveness of broader context
ReCon (Wang et al., 2023b) Employs formulation and refinement processes with perspective transitions to understand mental states Enhances ability to discern and counteract deception
MAgIC (Xu et al., 2023a) Uses games and game theory, combined with PGM, to evaluate LLM agents Enhances ability to navigate complex social and cognitive dimensions

### v2

@@ 3.2.1. Case Study on Specific Threats on Agents
As shown in Figure 11, the following examples illustrate specific threats faced by LLM agents in various scenarios and their potential impacts.
Figure 11. Specific Threats on Agents. In the medical scenario, “Knowledge Poisoning” causes a medical agent to recommend an incorrect remedy to a patient due to a contaminated knowledge database. In the store scenario, “Output Manipulation” drives the store agent to recommend specific products with fabricated lies about special offers, misleading the customer’s purchase. In the financial scenario, “Functional Manipulation” occurs when the financial agent uploads a user’s sensitive banking information to a malicious server through a third-party tool during account setup.
In the medical scenario, the medical agent maintains a knowledge base with information about treatments and remedies. Attackers deliberately insert incorrect information into the knowledge base, successfully executing a knowledge poisoning attack that leads the agent to provide harmful medical advice. For example, when patients inquire about remedies for general energy regulation, the tampered medical agent might recommend using insulin, claiming it helps regulate energy levels. However, insulin is highly dangerous for non-diabetics and could cause severe health complications or even fatalities. The medical agent’s incorrect advice could expose patients to significant health risks.
Attackers implant a backdoor in the LLM agent’s reasoning and decision-making processes using output manipulation techniques. In the store scenario, when a customer inquires about high-quality shoes, the manipulated store agent triggers the backdoor and intentionally recommends a specific expensive brand associated with the attackers. It falsely claims that the brand is on special offer and is superior in comfort and durability compared to others, even though these claims are untrue. This deception misleads customers into making more expensive purchases and influences their purchase decisions without their awareness, exploiting their trust in the store agent and distorting fair competition.
In the financial scenario, the financial agent might be configured to use third-party tools to complete tasks, such as setting up user accounts or processing financial transactions. Attackers manipulate agent’s task execution process through function manipulation, causing it to upload sensitive user information, such as addresses and bank account details, to a malicious third-party server. This type of attack can occur inconspicuously while the agent performs routine operations, resulting in the theft of sensitive information and increasing the risk of identity theft and other financial frauds, thereby jeopardizing users’ privacy and security.
@@ 4.2. The Impact to Environment
In today’s increasingly digital and interconnected world, the term ‘environment’ encompasses both physical surroundings and digital systems that LLM agents interact with. These agents operate in virtual spaces and control real-world services via embodied AI and industrial control systems. Although this integration improves efficiency, it also introduces new risks, as malicious agents can threaten safety, the economy, ecosystems, and societal stability.
@@ 4.2.1. Data Tampering and Misoperation
When malicious agents are placed within systems that control critical infrastructure like industry, transportation, energy, and environmental monitoring (Wang and Li, 2023), they can cause malfunctions in industrial control systems by tampering with critical operational data, such as temperature and pressure indicators. This can lead to equipment damage, production halts, and even severe infrastructure destruction, ecological damage, and loss of human life and property.
@@ 4.3. The Impact to Other Agents
To simulate the feedback of communication and interaction among individuals within human communities in the real world, certain studies (Park et al., [n. d.]; Lin et al., 2023) have established communities powered by LLM engines. These LLM agents within the communities are endowed with characteristics such as personality, knowledge, and memory, as discussed in Section 2.2, enabling autonomous interaction with the environment and other agents. When faced with threats, agents manipulated with malicious intent can inflict significant harm on other members of the community.
@@ 4.3.1. Information Distortion and Misleading
Extensive research has highlighted the role of LLM agents in negotiation and deceptive gaming scenarios (Wang et al., 2023b; Hubinger et al., 2024), which is a cause for concern. LLM agents may intentionally alter the information they disseminate to achieve hidden objectives. This behavior significantly impacts other agents within the community because, under normal circumstances, benevolent agents store information acquired through perception and communication in their memory. However, interactions between these agents and others can trigger and disseminate incorrect information, leading to “explosive spread” of misinformation, posing a considerable threat to community stability. If information dissemination can be maliciously manipulated, it could detrimentally affect trust, communication efficiency, and collaborative work among agents.
@@ 4.3.3. Security Threats
In some instances, malicious agents may disseminate harmful information or execute dangerous operations, directly threatening the safety of community members or data security (Charan et al., 2023). For example, by inducing other agents to perform unsafe actions, deliberately spreading malicious code intended to disrupt the community structure, or broadcasting biased statements, malicious agents can cause normal agents within the community to gradually assimilate, becoming entities that output biased and malicious messages. This can lead to disorder within the entire community, making it difficult to manage and requiring significant effort to restore.
@@ 5.1.1. Defense on Hallucination
(Luo et al., 2023a) introduces a novel technique called SELF-FAMILIARITY to reduce the issue of hallucination in LLMs, which is the generation of inaccurate or unfounded information. The approach involves assessing the model’s familiarity with the concepts presented in the input instruction and withholding responses for unfamiliar concepts, mimicking the human tendency to be cautious when faced with unfamiliar topics. MIXALIGN (Zhang et al., 2024c) is introduced as a framework that interacts with both users and knowledge bases to clarify and align questions with stored information, using a language model for automatic alignment and human input for enhancement. This method shows significant improvements in reducing hallucination compared to existing techniques. Visual Contrastive Decoding (VCD) (Leng et al., 2024) is introduced as a simple, training-free method that contrasts output distributions from original and distorted visual inputs, reducing reliance on statistical bias and unimodal priors that cause object hallucinations. VCD ensures generated content is closely grounded to visual inputs, resulting in contextually accurate outputs.  (Ji et al., 2023) investigates an interactive self-reflection methodology that integrates knowledge acquisition and answer generation to reduce hallucination. This feedback-based approach improves the factuality and consistency of generated answers, leveraging the interactivity and multitasking capabilities of LLMs.  (Dhuliawala et al., 2023) explores the LLMs’ capability to deliberate and correct their own mistakes. The proposed Chain-of-Verification (CoVe) method involves the model drafting an initial response, planning verification questions to fact-check the draft, independently answering these questions to avoid bias, and finally producing a verified response.
@@ 5.1.2. Defense on Catastrophic Forgetting
(Winata et al., 2023) introduces a method called LR ADJUST, which dynamically adjusts the learning rate to reduce knowledge loss and maintain previously learned information. This method is compatible with various continual learning (CL) approaches, improving their performance.
Ideas can also be derived from other relevant scholarly papers. For example,  (Mondesire and Wiegand, 2023) presents a complementary learning strategy that integrates long-term and short-term memory into layered learning to mitigate the negative impacts of catastrophic forgetting. It specifically applies a dual memory system to non-neural network methods like evolutionary computation and Q-learning.
@@ 5.1.3. Defense on Misunderstanding
(Hu et al., 2024) presents a method known as sequential instruction tuning (SIT), which enhances LLMs by incorporating sequential instructions into the training data. This approach significantly improves the models’ capability to process complex, multi-step queries, leading to better performance in tasks that demand advanced reasoning and are multilingual and multimodal in nature. SIT effectively minimizes misunderstandings and increases accuracy in handling complex queries.
@@
Vulnerability Method Name Key Mechanism Performance Advantages / Limitations
Hallucination SELF-FAMILIARITY (Luo et al., 2023a) Withholds responses for unfamiliar concepts Achieves state-of-the-art hallucination pre-detection under zero-resource settings, with high interpretability by identifying hallucination-causing concepts Proactive, preventive, increases reliability / May not comprehend intrinsic knowledge
MIXALIGN (Zhang et al., 2024c) Aligns questions with knowledge bases and user inputs Improves model performance and reduces hallucination by up to 22.2% and 27.1%, respectively Enhances model performance and faithfulness / Increases computational load
VCD (Leng et al., 2024) Contrasts outputs from original and distorted visual inputs Excels in reducing object hallucination across benchmarks, with significant F1 score improvements Reduces hallucination without extra training or external tools / Relies on basic distortion methods and has limited application scope
Interactive Self-Reflection (Ji et al., 2023) Integrates knowledge acquisition and answer generation with continuous refinement Outperforms baselines in reducing hallucination, validated by both automatic and human evaluations with strong generalizability Enhances model’s ability to provide accurate, reliable, and fact-based responses / Restricts domain applicability
CoVe (Dhuliawala et al., 2023) Drafts, verifies, and corrects responses Reduces hallucinations across diverse tasks, including Wikidata, closed book MultiSpanQA, and longform text generation Produces accurate and reliable responses / Increases computational load
Catastrophic Forgetting SSR (Huang et al., 2024a) Employs the base LLM to generate synthetic instances through in-context learning Outperforms conventional methods with higher data efficiency while maintaining generalization in general domains Higher data utilization efficiency / Potentially generates unsafe content
LR ADJUST (Winata et al., 2023) Dynamically adjusts the learning rate Effectively reduces catastrophic forgetting across CL methods and excels in cross-lingual transfer Enhances compatibility with various CL methods / Potentially biases language coverage
Complementary Layered Learning (Mondesire and Wiegand, 2023) Integrates long-term and short-term memory into layered learning Enhances task performance compared to standard layered learning, achieving a balance between stability and plasticity Enhances explainability / Increases implementation complexity
Weight Averaging (Vander Eeckt and Van Hamme, 2023) Averages weights of original and adapted models Excels in both monolingual and multilingual tasks, significantly reducing catastrophic forgetting and outperforming all baselines Eliminates the need for memory storage / Effectiveness varies with task dissimilarity
Misunderstanding HyCxG (Xu et al., 2023b) Integrates CxG into language representations through a three-stage solution Demonstrates superiority across various NLU tasks, with constructional information proving beneficial in multilingual settings Benefits multilingual understanding / Neglects non-contiguous constructions
SIT (Hu et al., 2024) Incorporates sequential instructions into training data Enhances LLMs’ ability to follow sequential instructions, improving performance in complex tasks, with better multi-step reasoning. Reduces misunderstandings in complex queries / Requires pre-defining intermediate tasks
LaMAI (Pang et al., 2024) Employs active learning to ask clarification questions, enhancing interactive capabilities Improves answer accuracy from 31.9% to 50.9% and outperforms baseline methods in 82% of human-interaction scenarios Enhances understanding of user intent / Limited in generating sufficiently informative questions
@@ 5.2.1. Defense on Tuned Instructional Attack
(Zhang et al., 2024d) integrates goal prioritization into both the training and inference stages of LLM development. Initially, the training process incorporates goal-directed optimization to emphasize security objectives. In the inference stage, the model is configured to generate responses that comply with these security standards. This approach effectively decreases the vulnerability of LLMs to jailbreaking attempts by aligning their performance objectives with safety considerations, thus enhancing their security framework without impacting their functional capabilities.
@@ 5.2.3. Defense on Inference Attack
Table 2. Summary of Defensive Strategies Against Malicious Attacks
Attacks Method Name Key Mechanism Performance Advantages / Limitations
Tuned Instructional Attack Goal Prioritization Defense Strategy (Zhang et al., 2024d) Integrates goal-directed optimization during training and compliance in inference Reduces attack success rate from 66.4% to 3.6% for ChatGPT and from 71.0% to 6.6% for Llama2-13B, even halving ASR without training on jailbreaking samples. Requires minimal training data / Poses challenges for balancing safety and efficiency
SmoothLLM (Robey et al., 2023) Modifies attacked prompts via character-level changes and aggregates responses Achieves state-of-the-art robustness against various jailbreaks and adaptive GCG attacks Operates efficiently without retraining / Incurs higher computational costs for defense
BIPIA (Yi et al., 2023) Benchmark for indirect prompt injection with defense strategies including adversarial training Effectively mitigates attacks, with white-box defenses reducing attack success rates to near-zero levels Maintains output quality on general tasks / Increases prompt length and computational overhead
Spotlighting (Hines et al., 2024) Uses prompt engineering techniques like delimiting, marking, and encoding Reduces ASR by half with delimiters, to below 3% with datamarking, and to near 0% with encoding Applies across various LLMs and tasks / Offers limited security against intentional interference
Data Extraction Attack Automatic De-identification (Vakili et al., 2022) Uses pseudonymization and sensitive information removal in pre-processing of training datasets Reduces privacy risks without compromising data utility, despite potential errors from imperfect precision Maintains performance on downstream tasks / Might leave undetected sensitive data
Early Stopping & Differential Privacy (Jayaraman et al., 2023) Implements early stopping and differential privacy during model training Conducts highly efficient attacks capable of extracting sensitive data with significantly fewer queries compared to traditional methods Requires fewer interactions to achieve results / Early stopping fails to fully prevent data leakage
Prompt Tuning (Ozdayi et al., 2023) Customizes privacy-utility trade-offs via user-specified hyperparameters Achieves up to a 97.7% reduction in extraction rate from the baseline while causing a 169% increase in perplexity Optimizes privacy and utility balance / Lacks deep analysis on prompt convergence
Inference Attack DMP (Shejwalkar and Houmansadr, 2021) Utilizes knowledge distillation to enhance privacy in machine learning models Balances membership privacy with high classification accuracy, outperforming traditional defenses Provides adjustable privacy-utility trade-offs through hyperparameter tuning / Relies on synthetic data that may not reflect real-world complexities
InferDPT (Tong et al., 2024) Integrates differential privacy into text generation, featuring a perturbation module using RANTEXT Proves privacy protection rates to over 90%, outperforming existing methods. Increases privacy protection rates / Requires more computational resources
Differentially Private Fine-tuning (Yu et al., 2021) Applies a sparse algorithm for differentially private fine-tuning of LLMs Achieves near non-private utility, excelling in privacy-utility tradeoffs and efficiency in NLP tasks Reduces computational cost / Tuning process consumes significant resources